OpenStudy (anonymous):

Compute the volume for the solid of revolution generated by rotating the region bounded by the curve y = x3 + 1, x-axis and the lines x = 3, y = 9 around the line x = 3.

OpenStudy (anonymous):

wat?

OpenStudy (anonymous):

I know how to compute the area of the region but what next?

OpenStudy (anonymous):

the curve is x^3 + 1 not x3 +1. sorry

OpenStudy (amistre64):

so we got something like this right?

OpenStudy (amistre64):

since we are rotating around x=3; we can use the shell method easily to find the volume

OpenStudy (amistre64):

(x+3)^3 +1 = 0 (x+3)^3 = -1 x+3 = -1 x = -4 ... our integration will be from -4 to 0 then

OpenStudy (amistre64):

\[2pi\int_{-4}^{0}x[f(x)]dx\] \[2pi\int_{-4}^{0}x[(x+3)^3+1]dx\] expand it and do the power rules

OpenStudy (amistre64):

check that y=9 part tho, 9 = (x)^3 +1 8 = (x)^3 2 = x when x = -1 ... so thats a part i dint account for :)

OpenStudy (amistre64):

2-3= -1; so integrate that first part from -4 to -1; and add the cylindar volume formed by a radius = 1 height = 9

OpenStudy (amistre64):